Sci IPT 20 Aug

Mangum, Wootten, Hills, Nyman, Lucas, Reid, Hales, Brogan, Hunter, Hibbard, Nikolic, Emerson

* ALMA Status Report R. Hills. See ALMA_Status_July08.pdf: for a recent set of slides. (These are more or less what was given at the URSI GA last week. I think you will have seen most of them before but there are some new pictures.) o Site: Installation of foundations continues. There are still several concerns here: the stiffness they provide; the design and positioning of the steel inserts that the antennas sit on; the practicality of the vaults where the power and fibres are connected. A new plan has been drawn up for converting the building at the OSF such that it provides what is needed. Unfortunately it will be well into next year before all the work can be done. Presumably we will be working in temporary arrangements until then. o Antennas: Testing of in MELCO #2 anntenna is complete and provisional acceptance (with the issue of the thermal behaviour still to be resolved) will occur as soon as the documentation is complete. The installation of the first ALMA FE and the associated electronics should take place in the next couple of weeks. More holography data is being taken on #1 to investigate the thermal problem. As noted above Vertex #1 and #2 are doing pointing tests. AEM #1 is undergoing drive tests at the factory. DTE reports holo on antenna Vx1 is below 11 microns. Probably managed to hit the asymptotic limit for the effort available now. Worst daytime measurement was 16 microns, 13-14 normally, still being written up with report in about a week. Change in focal length about 2x higher than expected from FEM seen with thermal change, still puzzling that one. This looks like a great antenna. REH: I thought the Focal length change with temp was 5x. DTE: That may be correct. REH: The surface rms associated with temp change exceeds FEM predictions also. JGM: focal length change from photogrammetry does not agree with that from holography, one needs to be cautious. We need to understand the details before discussing this further. o Front Ends: Delivery - the second front-end will be another "engineering model" - i.e. not fully tested - delivered from Taiwan. I believe that is now not expected to arrive until ~Nov 2008. Integration centers: Todd Hunter has agreed to take on the role of NA Front End scientist. There are still some aspects of the beam data that are not fully resolved. WVR - a partial CDR's on the IF section is taking place today. TH: Ade asked me to take on this role at 20% which will dominate my ALMA time for next few months. Attend meetings thru Aug then jump on. Perhaps LO RTM phase. o Back End: the second antenna system has been delivered and is full production activities cranking up. Plan is to deliver baseline system. Might work in parallel on alternative system which uses a different way to measure roundtrip phase, which may prove less likely to cause fringe skipping. The plan to have laser synthesizers 3 and 4 committed to the alternative plan not viable, we continue to work on this one. People involved really need to work on getting baseline done. o Correlator: some of the first quadrant of 64-input system has arrived in Chile. o Software: CDR6 completed. Overall external review planned for 17-23 Nov. o Systems: Revisit to the Systems Requirements planned - we need to discuss this. Laing on holiday, those involved not on line. This review would be to discuss if the reqs were captured along with flowdown. SE thought they needed more input from Sci side on how system would be used. I' flummoxed on what this really means. I will almost certainly need help; we'll have to put some effort into this. There is some reorganization of SE Whyborn is interim Systems Engineer; we will work with him. Schedule is a big concern at the moment. Schedule: A major effort is underway to stop the continuous slippage of the schedule that has been occurring. The steps are first to go through all deliverables with the managers responsible and agree a new set of realistic completion dates including modest but realistic contingency. The overall schedule will then be reassembled - which will I fear provide painful reading - and put under the control of a Schedule Control Board. Managers will have to request changes and explain why they have become necessary. This will occur between now and the end of October. This will, no doubt, result in some disappointments and may affect Early Science. * ASAC Matters. o Face-to-face meeting in Charlottesville after Massive Star Formation Workshop o Panel to prepare ALMA Development Proposal has had its first telecon

Science IPT

* Change Control Board items, other technical concerns (Hills) o Band 3 Receiver Noise Temperature Specification This is a request to relax the sensitivity requirements for this band (although the requirement for the outer edges of the band is tightened). The overall loss in sensitivity is not very large (perhaps 2 or 3 percent) but of course we don't want to give this up unless absolutely necessary. I have therefore put in a rather negative preliminary response. 41K in the center 48 in the outer parts. TH notes 50% meet 37K spec, others don't. Saini charged to rewrite CRE. Wants change in top-level specs. REH: Also this band takes 7% hit from goretex. TH: We could look at this as getting back more of the band. TH Lichtenberger sent 2 wafers, one unusable, one dirty contacts. Still wanted CRE as AL won't be able to redo chips soon. REH to decide if Board must be involved. TH to check wafers and warm optics. o 64-Antenna Correlator Specifications and Requirements Version B to Version C Change Request Four modes in Table 4 of the specification, ALMA-, were found to be impossible to implement in the detailed implementation phase due to a lack of resources in the correlator. These modes have been deleted from Table 4. Similar but slightly less sensitive modes are provided however. Note file AlternateModes?.doc in Lacasse's response please. o Band 7 cartridges: Cross-Polarisation Improvement This is (I trust!) just formal confirmation of the improvement already agreed. Long discussion on Configuration changes for the largest configuration. This is probably costly for the improvement in imaging we have been able to demonstrate. We can present this to the ASAC and listen to their advice but to push this through would take considerable effort and money. RR a few pad changes will not suffice. ATF: # ATF will operate thru 20 Dec, if at all possible. Lesser Sci IPT involvement, likely to just include a few experts. # Present goals: (1) Getting software working (2) Developing procedures needed for AIVC, and (3) Training personnel # Presently testing software version 5.1.1 which includes facilities for taking total power data and for using the ACA correlator. Total power modes are working well, and we are doing interferometry using control scripts written by Sci IPT members which will be tested and then turned into ALMA OT modes. # Hardware is in a fragile state and maintenance in Sept will be essential if operations are to continue. # The current schedule is below. Note that R. Laing will take over scheduling visiting science staff on Sept 1, and Debra will continue to schedule daily shifts at the ATF. During the current period foci of interest include further understanding of the behavior of the Line Length Corrector, identifying causes of latency, writing and exercising new control scripts, and reviewing the MS data more carefully to make sure the current version of the filler at the ATF is converting data properly.

Commissioning and Science Verification

* Plan. * Rainer Mauersberger and Robert Lucas have arrived in Chile and are already hard at work. * Antonio Hales will be starting work in Chile in October as an ALMA Commissioning Scientist. * NAOJ has appointed Dr. Tsuyoshi Sawada as ALMA Commissioning Scientist; he will probably arrive in Chile in October of this year. * Dr. Stuartt Corder has accepted a Jansky Postdoc; he will spend two years in Chile working with the Commissioning Team before moving to Charlottesville for the science portion of his appointment. * A request for additional manpower has cleared the Directors Council and should go to the Board on 25th Sept.

Jeff Mangum -- Calibration * Ancillary Calibration Devices: o Documents which describe the weather station and temperature profile technical specifications have been developed. o Initial purchase of (P,T,RH,Ws,Wdir) measurement devices has been made by Leonardo. o Specification of site infrastructural components (enclosure, tower, etc.) have been made and sent to Site IPT. * Refractive Bending and Delay Calculations: o Analysis complete. See wiki page for details. o Recent analysis of Yan refractive bending model suggests that it is limited (using sla_refro as benchmark). o Worked with Pat Wallace to implement improvements to sla_refro, including incorporation of Liebe atmospheric model. o Recommendation is to use sla_refro for refractive bending (note: currently used in M&C software in use at ATF). * ALMA Calibration Device (ACD): Several reviews (involving acronyms that I won't repeat) of ALMA Calibration Device (ACD) over past month point to both good and bad: o Good: Robotic arm tests suggest that positioning between bands will be as fast as 5-6 seconds, rather than the anticipated 9 seconds. o Bad: Frontend IPT unilaterally changed specification of solar filter to exclude use at Bands 1 and 2. Change not approved, and justification for change flawed. REH: Solar filter--I have an uneasy feeling there may be requirements we haven't matched. Al on solar flare calculation. Bastian is the local expert. REH will send him some questions.

Robert Lucas -- SSR * SSR meeting this month. (2008-08-13). Transition phase. Nyman will lead a subgroup on Ops, Peck a subgroup on CSV. Subroups will review testing and new requirements. Software acceptance in several times per year instances. * Testing of integrated system at the ATF, using data in ASDM format, version 5.0.3, by AIV and Science IPT. * Finalizing ASDM Finalization (version 1...) * Work starting n subarrays and on more flexible simulation in Software Test Environments (STE).

Disposition of existing equipment from site characterization * NRAO lends the 225 GHz tipper to ACT. ALMA may use a spare 183 GHz device for monitoring opacity. This will require a change in several Sci IPT planning documents, including the Site-Sci ICD. * ALMA continues to operate the NRAO radio interferometer (and the NRAO container and its power system). It is still working. Simon Radford downloaded the data a year ago and will now download the data taken during the year (there is room for about one year of data on the disk). The interferometer data reduction software was written by Mark Holdaway and is no longer operational (needs a Sun workstation). We may want to contract Mark to transfer the software to a LINUX system. Al will check on this. * There is a set of cameras on the NRAO container. We may consider to move them to the roof of the AOS TB in order to monitor and keep a record of the construction work. * Nyman will offer the two ESO weather stations and the broken ESO radio interferometer to the Argentinian/Brasilian site test group. * Wootten is investigating the status of the NRAO power system with the business office. * Radiosonde container will eventually migrate to OSF. * Lightning detector, seismometer and ESO atmospheric turbulence sounder status uncertain. Could use the prototypes for a general tipping radiometer.

OPT review--the design really isn't ready but we'll go ahead and may need to change course later. DTE: Documentation wasn't adequate. We left believing they are capable of building the device we required. We will make one and try it. WVR may also be late.

-- AlWootten - 20 Aug 2008
Topic revision: r1 - 2008-08-20, AlWootten
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback